ravencchatman:

Kamila plans to build a concrete block wall behind her house. The wall will be 12 feet long, 6 feet high, and 8 inches thick. Each concrete block measures 16 inches long by 8 inches wide by 8 inches deep. How many blocks will Kamila need to build the wall? Each concrete block has two holes through the middle of the block. The holes are 4 inches wide by 4.5 inches long. What is the volume of one concrete block? Each concrete block weighs 36.5 pounds. What is the total weight of the concrete wall? If each block is solid concrete, how much would one block weigh?

Extrinix:

I don't have time to answer it, but here are the equations: Question 1: Each concrete block/whole wall \(V=lwh\) Question 2: 2 holes: \(V=2(\pi 2^2(4.5)\) Volume of the wall with holes: \(V=8(8)(16)\) Volume of the wall without holes: \(V=8(8)(16) - \pi 2^2(4.5)\) Weight of blocks without holes: \(weight \times amount~of~blocks~without = total~weight\) Weight of blocks with holes: \(weight \times amount~of~blocks~with = total~weight\)

mhanifa:

Consider converting feet to inches when calculating volumes